OpenStudy (jazzyfa30):

use the given values to evaluate (if possible)all six trig functions. cot theta=-5 sin theta =(square root 26)/26

OpenStudy (jazzyfa30):

@johnweldon1993

OpenStudy (johnweldon1993):

Well you know that \[\large \sin(x) = \frac{\text{Opposite}}{\text{Hypotenuse}}\] so 

OpenStudy (johnweldon1993):

\[\large \cot(x) = \frac{1}{\tan(x)} = \frac{1}{\frac{sin(x)}{cos(x)}} = \frac{cos(x)}{sin(x)}\]

OpenStudy (johnweldon1993):

We also know that sin = √26/6 and that cot = -5 so \[\huge \frac{cos(x)}{\frac{\sqrt{26}}{26}} = -5\] \[\huge \frac{26cos(x)}{\sqrt{26}} = -5\] \[\large 26cos(x) = -5 \times \sqrt{26}\] \[\large cos(x) = \frac{-5\sqrt{26}}{26}\] Now from that we know \(\large cos(x) = \frac{\text{adjacent}}{\text{hypotenuse}}\) so that means our hypotenuse would be 26 (and you see above that it is...and our missing side is \(\large -5\sqrt{26}\)

OpenStudy (johnweldon1993):



OpenStudy (johnweldon1993):

And now you have your 3 sides to use for the trig functions...which I'm assuming you know